### v1

@@ Abstract
Bilayers (soft substrates coated with stiff films) are commonly found in nature with examples including skin tissue, vesicles, or organ membranes. They exhibit various types of instabilities when subjected to compression, depending on the contrast in material properties between the two components. We present wrinkling instabilities for 3D hyperelastic bilayer systems, including auxetics (materials with negative Poisson’s ratio), under uni-axial tension. In tension, a soft bilayer can experience large lateral contraction, and we find that with an adequate contrast in the Poisson ratios, compressive stresses may develop and generate wrinkles aligned with the tensile direction. We rely on an analytic modelling of the phenomenon, and validate it with a user-defined Python script with periodic boundary conditions and constitutive relation implementation in advanced Finite Element simulations. Our findings reveal that wrinkles are observed when the Poisson ratio of the substrate is greater than that of the film. As the two Poisson ratios converge to a common value, the critical stretch of instability shoots up rapidly, and the wrinkling disappears. We also confirm these results by asymptotic analysis. This wrinkling analysis has significant potential in controlling surface patterns of auxetic skin grafts and hydrogel organ patches under mechanical loads. Moreover, the asymptotic expressions in this work can be used under finite strain for buckling-based metrology applications.
keywords:
[…]
1 Introduction
@@
A survey of the literature shows that wrinkles appearing in tension are seldom studied beyond the linear-elastic framework [29], although they have been established experimentally under large strains [30, 15], see Fig. 1. In this work, we perform a linear buckling analysis of 3D hyperelastic thin stiff films on semi-infinite compliant substrates (including auxetics) in tension and large strains, in contrast to the numerous existing studies concerned with compression.
We first use the buckling analysis available in ABAQUS [31], which works up to a point, beyond which we have to rely on a semi-analytical approach with Mathematica [32]. We find that wrinkles may be generated in bilayers subjected to uni-axial tension. Main and original findings include (i) Deriving asymptotic expressions for critical stretch ratios and critical wavenumbers, valid in finite strains and useful for buckling-based metrology applications; (ii) Showing that the Poisson ratio of the substrate must be greater than that of the film for wrinkles to appear; (iii) Establishing that wrinkles disappear when the Poisson ratios of film and substrate converge to a common value, and that (iv) The wrinkling wavelength is high (low) when the Poisson ratios of film and substrate are close (far).
2 Results
2.1 3D bilayer system in tension
Figure 1: (a) Undeformed and (b) deformed configurations of a 3D bilayer system with periodic boundary conditions in the lateral directions. The system is infinite in the X- and Z-directions and is under uni-axial tension along the X-direction. The analysis shows that eventually, (b) wrinkles develop parallel to the X-direction, provided there is enough contrast between the constitutive parameters of the materials. (c) Image obtained using Scanning Electron Microscope (SEM) on a natural rubber substrate coated with a gold film, when elongated to 50% strain at room temperature [30]. An inset with an arrow shows the direction of elongation. Bright and dark bands represent the buckled gold film layer and the rubber showing through cracks in the stretched film, respectively. (d) Optical micrograph of a Polystyrene (PS) film layer on a silicon wafer with decreasing thickness from top to bottom is shown on the left panel. The corresponding wrinkles exhibit decreasing wavelengths on the right panel, when the PS films are attached to a Polydimethylsiloxane (PDMS) substrate and elongated to induced buckling [15].
We consider 3D bilayer systems with a thin stiff film perfectly bonded to a semi-infinite compliant substrate, see Fig. 1. Under a uniaxial tension applied along the X-direction, the bilayer is expected to eventually develop wrinkles aligned with that direction, as shown in Fig. 1(b), provided compressive stresses develop along the Y-direction because of a sufficient Poisson ratio contrast.
Figure 2: Comparison of wrinkling profiles in thin conventional and auxetic sheets. Top row: Thin acetate sheets under uni-axial tension, with (a) non-auxetic micro-structural patterns and (b) auxetic micro-structural patterns [33]. For conventional sheets, wrinkles develop at the center of the sheet. For auxetic sheets, they appear near the clamped edges. Bottom row: Buckling profiles (according to the magnitude of the displacement field) obtained with ABAQUS and the Blatz-Ko model Eq. 1. Here $\alpha=0.4$, $\mu=0.53$ GPa, and $\nu=0.38,-0.2$ in (c) and (d), respectively.
Here we use the Blatz-Ko strain energy function [34] for both materials:
$W=c_{1}\left(I_{1}-3+\frac{1}{\beta}\left(I_{3}^{-\beta}-1\right)\right)\\ +c_{2}\left(\frac{I_{2}}{I_{3}}-3+\frac{1}{\beta}\left(I_{3}^{\beta}-1\right)\right),$ (1)
@@ 2 Results
where $c_{1}\!=\!\alpha\frac{\mu}{2}$, $c_{2}\!=\!(1-\alpha)\frac{\mu}{2}$, $\beta=\frac{\nu}{1-2\nu}$, $\mathbf{F}$ is the deformation gradient, and $I_{1}=\text{tr}(\mathbf{FF}^{T})$, $I_{2}=I_{3}\text{tr}[(\mathbf{FF}^{T})^{-1}$], $I_{3}=\det(\mathbf{FF}^{T})$ are three strain invariants. Also, the material constants are the non-dimensional parameter $10<\alpha<$ 1, the initial shear modulus $\mu>0$, and the Poisson ratio $-1<\nu\leq 1/2$. We use the letters f and s for film and substrate, respectively. For the simulations, $\alpha_{f}=\alpha_{s}=0.4$ and $\mu_{f}/\mu_{s}=30$ (similar observations to those detailed in this study can be made for other values).
@@
Auxetic materials are highly compressible due to their negative Poisson ratio (thus far away from the incompressibility limit of a Poisson ratio equal to 1/2). Blatz-Ko strain energy functions with a negative Poisson ratio can thus be used to model auxetics, although with certain limitations [35]. For example, Ciambella and Saccomandi [36] use the Blatz-Ko strain energy to capture the experiments of Choi and Lakes [37] on auxetic materials. The model is also adequate to capture the experimental behaviors of membranes (no substrate) under tension, from auxetic to conventional Poisson ratios, see Fig. 2.
@@
Now, using the Blatz-Ko model for 3D bilayer systems under large uni-axial elongation, we find that wrinkles are observed only when the Poisson ratio of the film ($\nu_{f}$) is smaller than that of the substrate ($\nu_{s}$), as summarised in Fig. 3. This result recovers the work by Nikravesh et al. [38] conducted for linear-elastic bilayer systems.
Figure 3: Wrinkling condition according to the contrast between the Poisson ratios of film and substrate.
2.2 Incompressible substrate
Figure 4: Numerical simulations using Blatz-Ko material model (Eq. 1). Variations of the critical stretch of wrinkling $\lambda_{c}$ and corresponding critical wavenumber measure $k_{c}H$ with the Poisson ratio of one layer ($k_{c}$ is the critical wavenumber and $H$ is the film initial thickness). (a-b): The substrate is quasi-incompressible ($\nu_{s}=0.495$). (c-d): The film is highly auxetic ($\nu_{f}=-0.95$). Results from ABAQUS: black dots, results from Mathematica: solid line with square markers.
First, we assume that the substrate is quasi-incompressible ($\nu_{s}=0.495$) while for the film, $\nu_{f}\in(-0.95,0.495)$.
Here the Poisson ratio of the substrate is greater than that of film, and under uniaxial tension, the film expands more than the substrate along the Z-direction, resulting in compressive stresses and the formation of wrinkles parallel to the X-direction.
In Fig. 4(a-b), the values of the critical stretch of wrinkling $\lambda_{c}$ and of the corresponding critical wavenumber $k_{c}$ are plotted against $\nu_{f}$. From Fig. 4a, we observe that the critical stretch ratio values predicted from the semi-analytical analysis (Mathematica) and the FE buckling analysis ABAQUS match well, as long as $\nu_{f}\leq 0.35$. Beyond this value, ABAQUS ceases to predict wrinkles in the desired direction and gives negative eigenvalues, suggesting the load direction has to be reversed to obtain wrinkles, which is unphysical.
An auxetic film ($\nu_{f}<0$) expands in all directions under tension while the incompressible substrate contracts along the Z direction, leading to compressive stress in the film. Hence only low values of the critical stretch are required for the wrinkles to occur. As the Poisson ratio of the film increases and moves closer to that of the substrate, the values of the critical stretch shoot up sharply, supporting our findings in Fig. 3.
2.3 Highly auxetic film
Now we take the film to be highly auxetic, with Poisson ratio $\nu_{f}=-0.95$, while for the substrate, $\nu_{s}\in(-0.95,0.495]$.
@@ 2.3 Highly auxetic film
Fig. 4(c-d) show the variations of $\lambda_{c}$ and $k_{c}$ with $\nu_{s}$. Again, we find good agreement between analysis and simulations for a certain range, when $-0.7\leq\nu_{s}\leq 0.495$. For $\nu_{\textbf{s}}<-0.8$, ABAQUS stops providing meaningful predictions. In general, the buckling occurs early, except as $\nu_{s}\to\nu_{f}$ when it increases dramatically because in that limit, both film and substrate experience the same transverse contraction so that no compressive stress nor wrinkles develop. Similarly to Fig. 4b, Fig. 4d shows that the higher the Poisson ratio contrast is, the larger the wavelength of the wrinkles is.
@@
2.4 Asymptotic solution: Compressible neo–Hookean bilayer system
The following critical strain and wavenumber expressions:
@@ 2.4 Asymptotic solution: Compressible neo–Hookean bilayer system
have been derived by [30, 39, 38] for linearly elastic bilayers under uniaxial tension. These expressions are valid under the plane-strain approximation. Here, $h$ is the current thickness of the film, $\varepsilon_{c}$ is the critical strain, and $k_{c}$ is the critical wavenumber.
Figure 5: Numerical simulations for a bilayer made of compressible neo-Hookean materials ($\alpha_{s}=\alpha_{f}=1$ in Eq. 1). The substrate is quasi-incompressible ($\nu_{s}=0.495$). (a-b): Variations of the critical stretch of wrinkling $\lambda_{c}$ and corresponding critical wavenumber measure $k_{c}h$ ($k_{c}$: critical wavenumber, $h$: current film thickness) with the Poisson ratio of film layer $\nu_{f}$. ABAQUS: black dots, Mathematica: solid line with square markers, Asymptotic expressions (Eqs. 3 to 4): star markers, Linear-elastic expressions (Eq. 2): dash-dotted line.
However, we note that the plane-strain assumption is not valid for a bilayer system under uni-axial tension. Here, we follow Cai and Fu [40, 41] and assume that the shear modulus ratio $r={\mu_{s}}/{\mu_{f}}$ is of order $(k_{c}h)^{3}$ to derive asymptotic expansions for the critical strain and wavenumber (for simplicity, we take $\alpha_{s}=\alpha_{f}=1$ in Eq. 1). We find

### v2

@@ Abstract
Bilayers (soft substrates coated with stiff films) are commonly found in nature with examples including skin tissue, vesicles, or organ membranes. They exhibit various types of instabilities when subjected to compression, depending on the contrast in material properties between the two components. In this letter, we unravel the mechanisms behind wrinkling instabilities in auxetic bilayer systems under uni-axial tension. We find that in tension, a soft bilayer can experience large lateral contraction. With an adequate contrast in the Poisson ratios, compressive stresses may develop and generate wrinkles aligned with the tensile direction. We model analytically the onset of wrinkles and develop a new numerical method to validate our predictions with advanced Finite Element simulations in ABAQUS. Our findings reveal that wrinkles are observed when the Poisson ratio of the substrate is greater than that of the film. As the two Poisson ratios converge to a common value, the critical stretch of instability shoots up rapidly, and the wrinkling disappears. We also confirm these results by asymptotic analysis. Our proposed numerical method is versatile and can be used to predict wrinkling in bilayers where auxeticity is associated with microstructural patterns. Using inverse analysis, we design the film microstructures to achieve desired effective Poisson’s ratios and further validate the effective properties of these microstructures with the Finite Element code, FEAP. We show that the critical stretch ratio for buckling in auxetic structures with microstructural patterns is in strong agreement with the predictions of homogenized models. The new method proposed in this letter has a significant potential for controlling surface patterns of auxetic skin grafts and hydrogel organ patches under mechanical loads. Moreover, the asymptotic expressions in this work can be used under finite strain for buckling-based metrology applications.
preprint: APS/123-QED
@@
Figure 1: Left: Optical micrograph of a Polystyrene film on a silicon substrate taken from Stafford et al. [15]. After stretching the film wrinkles with different wavelength depending on the thickness of the film (decreasing from top to bottom insets). Right: Scanning Electron Microscope imaging of a natural rubber substrate coated with a gold film, when elongated (arrow) to 50% strain [19].
@@
A survey of the literature shows that wrinkles appearing in tension are seldom studied beyond the linear-elastic framework [30], although they have been established experimentally under large strains [19, 15], see Figure 1.
In this letter, we unravel the wrinkling instabilities of auxetic bilayers made of a thin stiff film bonded to a compliant substrate undergoing large tension. Under a uniaxial tension applied along the X-direction, such bilayers is expected to eventually develop wrinkles aligned with that direction, as shown in Figure 1 (right), provided compressive stresses develop along the Y-direction because of a sufficient Poisson ratio contrast. We discover that wrinkles are generated in bilayers subjected to uni-axial tension only when the Poisson ratio of the substrate is greater than that of the film. Moreover, we devise a new approach which uses a combination of semi-analytical and numerical methods to predict the wrinkling pattern of the bilayer in tension. This approach allows us to establish that wrinkles disappear when the Poisson ratios of film and substrate converge to a common value. Typically, the wavelength of wrinkles increases as the Poisson’s ratios of the film and substrate approach each other and decreases as they diverge. However, when the film and substrate are highly compressible $(\nu_{f}=-0.95\hskip 2.84526pt\text{and}\hskip 2.84526pt\nu_{s}<-0.7)$, this trend reverses, and the wavelength decreases, as illustrated in Figure 3(d).
We also derive asymptotic expressions for critical stretch ratios and critical wavenumbers. These new expressions are valid in finite strains and useful for buckling-based metrology applications. Finally, we employ an inverse design approach to achieve a desired effective Poisson’s ratio in the film by altering its microstructure. This method can be applied to the manufacturing of skin grafts, where the graft’s microstruture is chosen to avoid undesired wrinkling.
We start by implementing Finite Element simulations of a bilayer in tension in ABAQUS [31](see Supplementary material: Numerical model). We model the mechanical behaviour of each layer with the Blatz–Ko strain energy function [32] in the following form:
$W=c_{1}(I_{1}-3+\nicefrac{{(I_{3}^{-\beta}-1)}}{{\beta}})+c_{2}(\nicefrac{{I_{2}}}{{I_{3}}}-3+\nicefrac{{(I_{3}^{\beta}-1)}}{{\beta}}),$ (1)
@@
where $c_{1}\!=\!\alpha\frac{\mu}{2}$, $c_{2}\!=\!(1-\alpha)\frac{\mu}{2}$, $\beta=\frac{\nu}{1-2\nu}$, $\mathbf{F}$ is the deformation gradient and $I_{1}=\text{tr}(\mathbf{FF}^{T})$, $I_{2}=I_{3}\text{tr}[(\mathbf{FF}^{T})^{-1}$], $I_{3}=\det(\mathbf{FF}^{T})$ are the principal invariants of $\textbf{B}=\textbf{F}\textbf{F}^{\text{T}}$. The material constants are the non-dimensional parameter $10<\alpha<$ 1, the initial shear modulus $\mu>0$, and the Poisson ratio $-1<\nu\leq 1/2$.
@@
Auxetic materials are highly compressible due to their negative Poisson ratio (thus far away from the incompressibility limit of a Poisson ratio equal to 1/2). Blatz–Ko strain energy functions with a negative Poisson ratio can thus be used to model auxetics, although with certain limitations [33].
Figure 2: Thin acetate sheets (top) under uni-axial tension, with (a) non-auxetic micro-structural patterns and (b) auxetic micro-structural patterns [34]. For conventional sheets, wrinkles develop at the center of the sheet. For auxetic sheets, they appear near the clamped edges. Corresponding buckling profiles (bottom) obtained in ABAQUS with the Blatz–Ko model in Equation (1). Here $\alpha=0.4$, $\mu=0.53$ GPa, and $\nu=0.38,-0.2$ in (c) and (d), respectively.
For example, Ciambella and Saccomandi [35] use the Blatz–Ko strain energy to capture the experiments of Choi and Lakes [36] on auxetic materials. The model is also adequate to capture the experimental behaviors of membranes (no substrate) under tension, from auxetic to conventional Poisson ratios, see Figure 2. We thus implement a custom-made material subroutine in ABAQUS for the energy in Equation (1) and use it to model the mechanical behavior of the two layers. To induce the wrinkling pattern in the bi-layer under tension, we write a custom-made python script which implements periodic boundary conditions as shown in Figure 3(a) (see the section Finite Element Simulations in the Supplementary Material for further details on the numerical model).
Figure 3: (a) Our model of a bilayer system with periodic boundary conditions. The system is infinite in the X- and Z-directions and is stretched along the X-direction. Our numerical results show that wrinkles develop parallel to the X-direction, provided there is enough contrast between the mechanical parameters of the materials. (b) Wrinkling condition according to the contrast between the Poisson ratios of film and substrate. (c-d) Critical stretch $\lambda_{c}$ for wrinkling and corresponding non-dimensional critical wavenumber $k_{c}H$ plotted against the Poisson ratio of one layer ($k_{c}$ is the critical wavenumber and $H$ is the film initial thickness). In (c) the substrate is quasi-incompressible ($\nu_{s}=0.495$) and in (d) the film is highly auxetic ($\nu_{f}=-0.95$).
@@
Now, using the linear buckling analysis in ABAQUS, we find that the bilayer in tension develops wrinkles only when the Poisson ratio of the film ($\nu_{f}$) is smaller than that of the substrate ($\nu_{s}$), as summarised in Figure 3 (b). This result recovers the predictions obtained by Nikravesh et al. [37] for linear-elastic bilayer systems.
To predict the critical stretch and wavenumber of the wrinkling instability, we perform a linear stability analysis (see the section Semi-analytical Method in the Supplementary Material). We first calculate the base elastic solution. We then add a small incremental deformation to perturb the base solution and derive the associated incremental problem. Finally, we numerically solve the incremental problem in Mathematica [38] and find the critical stretch and wavenumber for the onset of wrinkling. To illustrate our results, we fix the Poisson ratio of the bilayer while varying that of the film, and vice versa. In Figure 3(c), we plot the critical stretch $\lambda_{c}$ and the corresponding wrinkling wavenumber $k_{c}$ against $\nu_{f}$ for a bilayer with a quasi-incompressible substrate. When the substrate is quasi-incompressible ($\nu_{s}\!=\!0.495$) and the film’s Poisson ratio ranges in $\nu_{f}\in(-0.95,0.495)$, we observe wrinkles. Here the Poisson ratio of the substrate is greater than that of film, and under uniaxial tension, the film expands more than the substrate along the Z-direction, resulting in compressive stresses and the formation of wrinkles parallel to the X-direction. We observe that the critical stretch ratio values predicted from the semi-analytical analysis (Mathematica) and the FE buckling analysis (ABAQUS) match well, as long as $\nu_{f}\leq 0.35$. Beyond this value, ABAQUS ceases to predict wrinkles in the desired direction and gives negative eigenvalues, suggesting the load direction has to be reversed to obtain wrinkles, which is unphysical. An auxetic film ($\nu_{f}<0$) expands in all directions under tension while the incompressible substrate contracts along the Z direction, leading to a compressive stress in the film. Hence only low values of the critical stretch are required for the wrinkles to occur. As the Poisson ratio of the film increases and moves closer to that of the substrate, the values of the critical stretch shoot up sharply, supporting our findings in Figure 3(b).
Figure 4: Comparison between our asymptotic results (red star markers), linear asymptotic results (dashed lines), numerical simulations (black squares) and linear stability analysis (empty blue squares). Both substrate and film are neo–Hookean ($\alpha_{s}\!=\!\alpha_{f}\!=\!1$ in Equation (1)) and the substrate is quasi-incompressible ($\nu_{s}\!=\!0.495$). (a-b): Critical stretch $\lambda_{c}$ and corresponding critical wavenumber $k_{c}h$ against the Poisson ratio of film $\nu_{f}$. The asymptotic equations for the linear case are reported in Equation 2.
Figure 5: (a) Influence of microstructural geometry on the effective Poisson’s ratio ($\nu_{f}^{\text{eff}}$) in auxetic films with orthogonal-oval and rectangular-void patterns. Films with desired effective Poisson’s ratios are obtained using inverse analysis. (b) Comparison of undeformed and deformed configurations in bilayers with distinct auxetic microstructures, each designed to achieve specific Poisson’s ratio in the small strain regime. Film and substrate base materials exhibit compressible ($\nu_{f}=0.1$) and nearly incompressible ($\nu_{f}=0.495$) properties, respectively. Periodic boundary conditions are applied on the lateral faces of the domain along X and Z and the system is subjected to uni-axial tension along X. We observe that the wrinkles are generated parallel to the direction of tension. The bottom of (b) shows the percentage relative error in critical stretch between homogenized bilayers (Figure 3(c)) and those with microstructural patterns, obtained in ABAQUS with the Blatz–Ko model.
@@
We also observe wrinkles when we take the film to be highly auxetic (with Poisson ratio $\nu_{f}=-0.95$) and the substrate Poisson ratio varies in $\nu_{s}\in(-0.95,0.495]$. Figure 3(d) shows how $\lambda_{c}$ and $k_{c}$ vary with $\nu_{s}$ for a bilayer with a highly auxetic film. Again, we find good agreement between theory and simulations for a certain range, when $-0.7\leq\nu_{s}\leq 0.495$. For $\nu_{\textbf{s}}<-0.8$, ABAQUS stops providing meaningful predictions. In both scenarios, wrinkles occur at low stretching values ($\lambda_{c}\!<\!1.5$), except as $\nu_{s}\!\to\!\nu_{f}$. In this limit, the critical stretch increases dramatically because both film and substrate experience the same transverse contraction so that no compressive stress nor wrinkles develop. Similarly to Figure 3(c), Figure 3(d) shows that the higher the Poisson ratio contrast is, the larger the wavelength of the wrinkles.
@@
Analytical expressions for bilayers under uni-axial tension have historically been derived only for linear elastic materials, as seen in the works of Volynskii et al. [19], Chung et al. [39], and Nikravesh et al. [37]. The critical strain and wavenumber expressions are given by:
@@
Here, $h$ is the current thickness of the film. These expressions, valid under the plane-strain approximation and at low strains ($\ll 10\%$), have been applied in buckling-based metrology to determine the Young’s modulus of the film [15].
However, for a bilayer system under uni-axial tension the plane-strain assumption is not valid. To address this, we employ asymptotic analysis to derive the expressions for $\lambda_{c}$ and $k_{c}h$ based on the material properties of film and substrate layers, in finite strain deformation. To derive universal asymptotic expansions in large deformations for the critical strain and wavenumber, we follow Cai and Fu [40, 41]. For simplicity, we take $\alpha_{s}\!=\!\alpha_{f}\!=1$ in Equation (1) and assume that the shear modulus ratio $r={\mu_{s}}/{\mu_{f}}$ is of order $(k_{c}h)^{3}$. We find the following asymptotic expressions: